An engineer has designed a valve that will regulate water pressure on an automobile engine. The valve was tested on 200 engines and the mean pressure was 4.5 lbs/square inch. Assume the standard deviation is known to be 0.7. If the valve was designed to produce a mean pressure of 4.4 lbs/square inch, is there sufficient evidence at the 0.01 level that the valve does not perform to the specifications?

State the null and alternative hypotheses for the above scenario.
Answer #1

Solution:

This is a two tailed test.

The null and alternative hypothesis is,

Ho: 4.4

Ha: 4.4

The test statistics,

Z =( - )/ (/n)

= ( 4.5 - 4.4 ) ( 0.7 / 200)

= 2.02

P-value = 2 * P(Z > z )

= 2 * ( 1 - P(Z < 2.02) )

= 2 * 0.0217

= 0.0434

The p-value is p = 0.0434, and since p = 0.0434 < 0.1, it is concluded that the null hypothesis is not rejected.